What is 2 over 3 - 1 over 3 over 3 over 4
Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

2/3-1/3/(3/4)
2/3-1/3*4/3
2/3-4/9
6/9-4/9=2/9 ANS.

Answer by rfer(16322)

2/3-(1/3)/(3/4)
2/3-4/9
6/9-4/9=2/9